The width of a rectangle is 6km less than twice its length. If its area is 216km, find the dimensions of the rectangle.
Here's what I came up with:
2x-6=216
2x-6+6=216+6
2x=222
x=111
I'm not sure about this. word problems stump me sometimes.

Answer by scott8148(6628)

let x=length, so 2x-6=width ___ (x)(2x-6)=216 ___ 2x^2-6x=216 ___ x^2-3x-108=0

(x-12)(x+9)=0